APPLICATION OF THE ’MLR’ DIRECT POTENTIAL FITTING (DPF) METHOD FOR THE X^{1}Σ_{g}^{+} STATE OF Cs_{2}

\underline{JOHN A. COXON}, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia B3H 4J3, Canada; PHOTOS G. HAJIGEORGIOU, Department of Life and Health Sciences, University of Nicosia, 46 Makedonitissas Avenue, P.O Box 24005, 1700 Nicosia, Cyprus.

A set of 16544 A^{1}Σ_{u}^{+} - X^{1}Σ_{g}^{+} fluorescence line positions^{a} with a measurement precision of 0.001 cm^{−1}, and which samples ground state levels up to υ′′ = 135 lying about 28 cm^{−1} below the dissociation limit, has been employed in least-squares fits using the MLR direct potential fitting approach ^{b}. As well as the extended MLR model employed recently for Li_{2}, a further extension employing three y_{i}-variables leads to an optimum model for Cs_{2} that ensures a realistic extrapolation into the long-range region in accord with the theoretical dispersion constants.

aC. Amiot and O. Dulieu, J. Chem. Phys. 117, 5155 (2002).

bR. J. Le Roy and R. D. E. Henderson, Mol. Phys. 105, 663 (2007).